OpenStudy (anonymous):

k(x)=x+7 k(0)=x+7=7 k(-10)=x+7=-3 k(-19)=x+7=26 k(7)=x+7=0 k(d+13)=x+7 Are these right?

OpenStudy (anonymous):

k(-19)=x+7=26 this one is wrong -19+7=-12 k(7)=x+7=0, 7+7=14 k(d+13)=x+7 d+13+7= d+20 the rest you right